# Triple-BERT: Do We Really Need MARL for Order Dispatch on Ride-Sharing Platforms?

### v1

@@ E.4 Additional Experiment Result
Lastly, we note that our method and the approach by Enders et al. [8] exhibit higher confirmation times. This may be attributed to both methods having an explicit rejection action (i.e., choosing no order), unlike the other methods. While this mechanism can lead to higher confirmation times, it also enables the model to discard negative profit orders and reserve some orders for currently unavailable workers.
@@
Method Reward Service Rate Delivery Time Detour Time Pickup Time Confirmation Time DeepPool [1] 12723.85 0.91 11.53 2.47 7.77 0.06 BMG-Q [16] 13036.29 0.92 10.57 1.90 7.61 0.10 HIVES [12] 12365.11 0.89 11.04 2.28 7.99 0.03 Enders et al. [8] 12041.62 0.90 12.28 2.90 7.94 0.80 CEVD [3] 13157.96 0.94 11.36 2.31 7.37 0.06 Triple-BERT 14730.48 0.98 11.53 2.52 5.73 0.13 w/o stage 1 10665.02 0.87 11.92 2.72 9.36 0.68 w/o normalization 10839.33 0.88 12.50 2.85 9.10 0.24
@@ E.5 Expanded Generalization and Extensibility Experiment
Scenario DeepPool [1] BMG-Q [16] HIVES [12] Enders et al. [8] CEVD [3] Triple-BERT 7.16 (4,451 orders) 13,473 14,121 12,070 12,142 14,226 16,831 7.17 (4,125 orders) 13,204 13,424 12,232 12,208 14,145 16,145 7.18 (3,635 orders) 12,679 13,067 12,397 12,268 13,336 14,819
@@
Driver Amount DeepPool [1] BMG-Q [16] HIVES [12] Enders et al. [8] CEVD [3] Triple-BERT 1,500 21,090 22,333 19,587 19,546 23,092 27,458 2,000 25,316 25,650 25,713 25,394 26,207 28,273
At last, to prove the practical application potential of our method, we test the computation time with driver amounts ranging from 1,000 to 2,000 and order amounts from 300 to 500 at a single time step, shown in Table 6. (In our real-world dataset, we typically observe that the order amount does not exceed 200 at any single step.) The results indicate that the decision time remains consistently under 0.2 seconds across all scenarios. Additionally, the order and driver counts have minimal impact on computation time. This suggests that, for the current simulation, most of the processing time is allocated to the simulator’s operations rather than to decision computation cost.
@@ E.6 Expanded Ablation Study
To illustrate the efficiency of the encoder, we compare our methods and others when using the attention-based encoder and vanilla MLP-based encoder. The result is shown in Table 8. The results indicate that the ARL-based encoder significantly enhances the performance of our Triple-BERT alongside independent MARL methods such as DeepPool and BMG-Q. However, this improvement does not extend to CTDE and centralized MARL approaches like HIVES, Enders et al., and CEVD. The underlying reason is that CTDE and centralized MARL methods in ride-sharing primarily suffer from the curse of dimensionality in the critic network, an issue that cannot be mitigated by a more powerful feature extractor. We believe that the improvements observed in DeepPool, BMG-Q, and our Triple-BERT effectively enhance the efficiency of our designed encoder. Furthermore, even with the ARL-based encoder, our Triple-BERT consistently outperforms all previous methods, underscoring the superiority of our approach. Additionally, even with the simple MLP-based encoder, Triple-BERT still outperforms previous MARL methods, illustrating the robustness of our centralized SARL method.
@@
Encoder DeepPool [1] BMG-Q [16] HIVES [12] Enders et al. [8] CEVD [3] Triple-BERT Vanilla MLP 13,332 13,539 13,126 12,670 13,746 14,967 ARL-based 14,570 13,879 13,095 12,706 13,724 15,388
@@ Appendix F Related Work
The order assignment has become an essential task in on-demand transportation platforms like ride-hailing, ride-sharing, and food delivery. Methods can be mainly divided into two types: model-based (optimization-based) and RL-based. Most model-based methods suffer from high computation time for optimization and struggle to capture future uncertainty and dynamics, making it hard to explore the global optimal solution. In contrast, RL-based methods, especially Deep RL, can make fast decisions using neural networks and learn environmental uncertainties through self-interaction.
In the early stages, most RL-based works were inspired by [2], following a paradigm that first estimates the Q-value by estimating the V-value of the next state and combining it with the reward, then using bipartite matching to maximize the global Q-value [36; 22; 24; 9]. However, the application scenario of such methods is very limited; when the reward function is comprehensive, the computation time becomes unacceptable. The Q-value computation approach is somewhat complex. Works like [16; 15] address this problem successfully by directly using DQN to estimate the Q-value of each potential state-action pair. Furthermore, some works follow this paradigm but use policy gradient-based methods for optimization [33; 38].
However, a shortcoming of the above works is that they depend on the independent assumption, which means each driver is viewed as an agent while others are seen as part of the environment. Even though the goal of the order assignment task is to maximize the global reward, this independent paradigm trains each agent to maximize their own reward, which can lead to unnecessary competition among agents and further diminish the global reward. Some methods [8; 12; 3] aim to address this problem by introducing a centralized critic, realizing the CTDE paradigm. However, without a carefully designed network, the large number of agents can lead to the curse of dimensionality, hindering efficient training.

### v2

@@ E.4 Additional Experiment Result
Lastly, we note that our method and the approach by Enders et al. [enders2023hybrid] exhibit higher confirmation times. This may be attributed to both methods having an explicit rejection action (i.e., choosing no order), unlike the other methods. While this mechanism can lead to higher confirmation times, it also enables the model to discard negative profit orders and reserve some orders for currently unavailable workers.
@@
Method Reward Service Rate Delivery Time Detour Time Pickup Time Confirmation Time DeepPool [al2019deeppool] 12723.85 0.91 11.53 2.47 7.77 0.06 BMG-Q [hu2025bmg] 13036.29 0.92 10.57 1.90 7.61 0.10 HIVES [hao2022hierarchical] 12365.11 0.89 11.04 2.28 7.99 0.03 Enders et al. [enders2023hybrid] 12041.62 0.90 12.28 2.90 7.94 0.80 CEVD [bose2023sustainable] 13157.96 0.94 11.36 2.31 7.37 0.06 Triple-BERT 14730.48 0.98 11.53 2.52 5.73 0.13 w/o stage 1 10665.02 0.87 11.92 2.72 9.36 0.68 w/o normalization 10839.33 0.88 12.50 2.85 9.10 0.24
@@ E.5 Expanded Generalization and Extensibility Experiment
Scenario DeepPool [al2019deeppool] BMG-Q [hu2025bmg] HIVES [hao2022hierarchical] Enders et al. [enders2023hybrid] CEVD [bose2023sustainable] Triple-BERT 7.16 (4,451 orders) 13,473 14,121 12,070 12,142 14,226 16,831 7.17 (4,125 orders) 13,204 13,424 12,232 12,208 14,145 16,145 7.18 (3,635 orders) 12,679 13,067 12,397 12,268 13,336 14,819
@@
Driver Amount DeepPool [al2019deeppool] BMG-Q [hu2025bmg] HIVES [hao2022hierarchical] Enders et al. [enders2023hybrid] CEVD [bose2023sustainable] Triple-BERT 1,500 21,090 22,333 19,587 19,546 23,092 27,458 2,000 25,316 25,650 25,713 25,394 26,207 28,273
Additionally, to prove the practical application potential of our method, we test the computation time with driver amounts ranging from 1,000 to 2,000 and order amounts from 300 to 500 at a single time step, shown in Table 7. (In our real-world dataset, we typically observe that the order amount does not exceed 200 at any single step.) The results indicate that the decision time remains consistently under 0.2 seconds across all scenarios. Additionally, the order and driver counts have minimal impact on computation time. This suggests that, for the current simulation, most of the processing time is allocated to the simulator’s operations rather than to decision computation cost.
@@
To further demonstrate the generalizability of our method, we conducted an expanded experiment using High Volume For-Hire Vehicle (FHV) trip data from Queens, New York City [TLCData]. We chose not to continue with the yellow taxi data, as its primary operational area is Manhattan. Unlike the capital-intensive region of Manhattan, Queens has a significantly lower trip volume, despite its area being about five times larger. Consequently, the data distribution in Queens presents a markedly different challenge: while the number of orders decreases, the distances between them tend to increase, leading to greater difficulties for ride-hailing services. To adapt to this scenario, we set the driver count to 500 while keeping all other settings consistent.
The detailed experimental results are presented in Table 8, using data from 19:00 to 19:30 on July 17, 2024, which includes 2,024 valid orders. We observe that our Triple-BERT maintains its SOTA performance, primarily optimizing the assignment and improving rewards by serving more orders and reducing pickup times. However, this inevitably leads to a slight increase in delivery and detour times due to the bundling of more orders.
Additionally, compared to the results in Manhattan, as shown in Table 4, we noted that the average pickup time in Queens is significantly longer. While the decrease in the number of drivers may contribute to this, it is not the main factor, as the order volume has also declined significantly. The primary reason lies in the larger area and more dispersed order distribution in Queens. This leads to a substantial penalty in the pickup time term of the reward function, resulting in minimal differences in the rewards among different policies. Such challenging scenarios further hinder the efficient exploration and learning of centralized MARL methods like HIVES and CEVD.
Table 8: Performance of Different Methods in Queens, New York City [TLCData]
Method Reward Service Rate Delivery Time Detour Time Pickup Time Confirmation Time DeepPool [al2019deeppool] 5222.85 0.64 11.24 1.84 12.30 0.21 BMG-Q [hu2025bmg] 5362.00 0.66 9.63 1.08 12.98 0.27 HIVES [hao2022hierarchical] 3560.80 0.60 8.30 0.36 14.67 0.41 Enders et al. [enders2023hybrid] 4543.68 0.61 10.41 0.85 13.39 2.25 CEVD [bose2023sustainable] 4388.83 0.62 11.61 1.33 13.74 0.29 Triple-BERT 5577.83 0.72 9.07 0.90 11.32 0.23
@@ E.6 Expanded Ablation Study
To illustrate the efficiency of the encoder, we compare our methods and others when using the attention-based encoder and vanilla MLP-based encoder. The result is shown in Table 10. The results indicate that the ARL-based encoder significantly enhances the performance of our Triple-BERT alongside independent MARL methods such as DeepPool and BMG-Q. However, this improvement does not extend to CTDE and centralized MARL approaches like HIVES, Enders et al., and CEVD. The underlying reason is that CTDE and centralized MARL methods in ride-sharing primarily suffer from the CoD in the critic network, an issue that cannot be mitigated by a more powerful feature extractor. We believe that the improvements observed in DeepPool, BMG-Q, and our Triple-BERT effectively enhance the efficiency of our designed encoder. Furthermore, even with the ARL-based encoder, our Triple-BERT consistently outperforms all previous methods, underscoring the superiority of our approach. Additionally, even with the simple MLP-based encoder, Triple-BERT still outperforms previous MARL methods, illustrating the robustness of our centralized SARL method.
@@
Encoder DeepPool [al2019deeppool] BMG-Q [hu2025bmg] HIVES [hao2022hierarchical] Enders et al. [enders2023hybrid] CEVD [bose2023sustainable] Triple-BERT Vanilla MLP 13,332 13,539 13,126 12,670 13,746 14,967 ARL-based 14,570 13,879 13,095 12,706 13,724 15,388
@@ Appendix F Related Work
F.1 Order Dispatch in Ride-Sharing Task
Order dispatch methods in ride-sharing can be primarily categorized into model-based and RL-based approaches. Model-based methods often rely on early assumptions that all order information is known in advance [yan2011model] or neglect potential future dynamics, resulting in impractical or myopic solutions. Later research focused on modeling and capturing environmental uncertainties for practical applications [alonso2017demand]. However, accurately characterizing these complexities in the ride-sharing market remains a challenging task.
In contrast, model-free RL methods free researchers from these constraints, allowing agents to interact with and learn from the environment independently. Given the vast action and observation spaces, most studies adopt a MARL paradigm (some referring to their methods as decentralized RL) to effectively manage these challenges [qin2025reinforcement]. Similar to standard MARL, the methods adapted for ride-sharing can be divided into Decentralized Execution (DTDE, also known as independent MARL), Centralized Training with Decentralized Execution (CTDE), and Centralized Training with Centralized Execution (CTCE) [jin2025comprehensive]. DTDE methods, while widely used in early studies [al2019deeppool; feng2022coordinating], suffer from low cooperation since each agent perceives others merely as part of the environment, often leading to instability during training. Hu et al. [hu2025bmg] introduced the GAT [velickovic2017graph] to enable each agent to consider the information of its neighbors, thereby reducing overestimation issues and improving system performance. Other studies have shifted towards CTDE and CTCE paradigms to promote effective cooperation. For instance, Enders et al. [enders2023hybrid] proposed a delayed matching method based on MASAC [haarnoja2018soft], allowing agents to decide whether to accept assigned orders at each step. Bose et al. [bose2023sustainable] developed a VD-based method [sunehag2018value] that utilizes a global reward to foster cooperation, although it faces the challenge of lazy agents. Furthermore, Hao et al. [hao2022hierarchical] applied QMIX [rashid2020monotonic], while Hoppe et al. [hoppe2024global] employed COMA [foerster2018counterfactual] alongside a mix of global and local rewards to address existing issues. However, many of these centralized methods require a centralized critic, reintroducing the challenge of CoD. To address this, Li et al. [li2019efficient] proposed a Mean Field MARL framework [yang2018mean], and Zhao et al. [zhao2025one] developed the GRPO [shao2024deepseekmath] and OSPO methods based on the homogeneous properties among agents, although these assumptions are not always met in practice.
Additionally, several studies began to explore more practical scenarios that integrate order dispatch with other tasks, such as repositioning, price setting, and multi-modal transportation [^5] To avoid confusion, we would like to clarify that the concept of multi-modal in transportation differs from that in computer science. In transportation, “multi-modal” refers to a travel model that utilizes multiple vehicles, such as taxis, buses, and subways, regardless of the model input.. For example, Zhang et al. [zhang2023future] considered the joint optimization of order dispatch and price setting, which can significantly influence customer demand. Similarly, Haliem et al. [haliem2021distributed] and Ge et al. [ge2025marl] examined repositioning strategies based on these factors. Hu et al. [hu2025coordinating; hu2025real] analyzed order assignment in joint delivery scenarios involving cars, subways, and Unmanned Aerial Vehicles (UAVs), while other works, such as Singh et al. [singh2021distributed], have studied multi-hop transportation. Furthermore, some studies have examined fairness considerations. For instance, Zhang et al. [zhang2023mutual] introduced mutual information in the reward function to address challenges related to unusual order distributions and improve platform income. Zhao et al. [zhao2025impacts] investigated algorithmic discrimination by considering joint order assignments and payment settings for workers.
Overall, current MARL-based ride-sharing methods face a fundamental dilemma: centralized training paradigms foster cooperation at the cost of exacerbating the CoD, whereas decentralized approaches mitigate CoD but suffer from non-stationarity and poor coordination, ultimately limiting system performance. To navigate this dilemma, our paper offers a novel perspective by directly utilizing a centralized SARL framework to leverage complete global information, where we then develop efficient techniques, such as action decomposition, to directly address the ensuing CoD challenge.
F.2 Cooperative Multi-Agent Reinforcement Learning
Cooperative MARL has a wide range of applications, including power control, robot fleet management, and ride-sharing tasks [yuan2023survey]. According to a recent survey [oroojlooy2023review], these methods can be categorized into five types: independent learning, centralized critic, value decomposition, consensus, and communication.
Early-stage works primarily focused on independent learners, which represent the simplest realization of MARL. By viewing each agent as an independent entity and others as part of the environment, standard SARL methods can be easily transferred to MARL scenarios. However, this approach can lead to unstable environments, as the policies of other agents continuously change. Additionally, it often results in local optima, as each agent aims to maximize its own reward, neglecting the cooperation necessary to optimize global reward. Despite efforts such as Hysteretic Q-Learning [matignon2007hysteretic], challenges remain in large-scale environments and sparse reward situations.
To overcome these challenges, the CTDE and CTCE paradigms have been widely explored in centralized critic and value decomposition MARL methods. In the case of centralized critics, early research focused on adapting actor-critic methods like PPO, SAC, and DDPG by replacing the critic with a centralized one. During training, the critic receives input from all agents, thus addressing the instability problem, although it is not needed during execution, allowing actors to function independently. To further enhance this paradigm, attention-based critics were introduced to capture the relationships among agents [mao2018modelling], inspired by the Transformer architecture [vaswani2017attention]. Foerster et al. [foerster2018counterfactual] proposed COMA to resolve the credit assignment problem by introducing a counterfactual baseline, which underwent further improvements in subsequent research [li2021shapley]. Additionally, value decomposition methods aim to effectively distribute global rewards among agents, enabling them to optimize the overall reward rather than just their individual rewards. Starting from VD [sunehag2018value], which struggles with lazy agents due to the simple addition of Q-values, many efforts have been made to enhance the representation capacity of functions that combine global and individual Q-values. Notable examples include QMIX [rashid2020monotonic], VDAC [su2021value], and QTRAN [son2019qtran]. However, most of these methods encounter the challenge of CoD when the number of agents increases, particularly in scenarios like our ride-sharing task, where the number of agents can exceed hundreds or even thousands.
Consensus and communication methods, developed later, strive to find a balance between low cooperation and the CoD challenge. Here, agents only exchange information with their neighbors or selected agents. Consensus-based methods utilize sparse communication to achieve policy consensus among agents, often with convergence guarantees under linear approximators [varshavskaya2009efficient]. However, many of these methods often rely on multiple rounds of communication, leading to practical challenges in the ride-sharing task, which has high demands for real-time performance. In contrast, communication-based methods focus on designing efficient mechanisms for determining what information to send and to whom. For instance, Sukhbaatar et al. [sukhbaatar2016learning] proposed CommNet, which broadcasts each agent’s hidden features derived from their observations. However, similar to Mean Field approaches [yang2018mean] in MARL, this method considers only the average influence of others, neglecting the specific relationships among agents. Subsequent attention-based methods were introduced to evaluate the significance of information from different agents [das2019tarmac; li2024context]. However, the training difficulties associated with communication-based methods are relatively high, particularly as early-stage communication often yields little meaningful information. Furthermore, many communication methods face a trade-off between cooperation performance, communication costs, and the content of communication—issues closely related to CoD. Even recent methods focusing on efficient communication mechanism design [wang2023ac2c] still encounter challenges in our ride-sharing scenario with extremely high agent volumes. For example, [liexponential] proposed a novel communication protocol based on exponential graphs, guaranteeing global message exchange within $\lceil\log_{2}(N-1)\rceil$ steps. However, this number can reach 10 in our experiment with 1,000 agents, which is relatively high compared to the total time step horizon of 30. Additionally, communication-based methods represent an ongoing challenge, as determining what content to communicate and to whom poses risks associated with credit assignment. Despite these advancements, most consensus and communication methods still encounter challenges such as credit assignment and a lack of convergence theory. In contrast, the SARL-based method circumvents these problems by directly capturing global information. Given that ride-sharing is fundamentally a centralized decision task, we consider the SARL-based method to be a promising solution that overcomes the limitations of previous MARL-based approaches.